OpenStudy (anonymous):

Unit circle tan(-600)? I know the answer to the question but I am confused, you can simplify tan(-600) by +360 and +180 to get tan(-60) but without doing that cant tan(-600) also be tan(120)? Basically whats a surefire way of doing large negative values (over 360)

OpenStudy (freckles):

you can do either one tan(-600) is tan(-60) or tan(120)

OpenStudy (alekos):

tan(-600) = tan(120) = - tan(60)

OpenStudy (anonymous):

Also like with csc(-270) that would be obviously be easily solved because it is equal to tan(90) so is it also allowed to add 180 to -270 and get -90 degrees? csc(-90) would give the same answer as csc(90) (undefined) soo.. is that a coincendence? or can you only subtract increments of 360/180 when the degree is over 360?

OpenStudy (freckles):

sin(-270)=1 so csc(-270)=1 but tan(90)=1/0 which is undefined so I don't see how you got csc(-270) is same as tan(90)

OpenStudy (freckles):

also csc(-90) isn't the same as csc(90) csc(-90)=-1 and csc(90)=1 since sin(-90)=-1 and sin(90)=1

OpenStudy (anonymous):

oh man sorry i meant sec and i think i understand now now under 360, I should solve as always and not add or subtract increments and while over increments of 180 or 360 that is when I should do that.